OpenStudy (anonymous):

The equation of the line the point (-1,3) and perpendicular to the line x+2y=-6 is

OpenStudy (anonymous):

what is the slope of x+2y=-6?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

No...put it in slope intercept form x+2y=-6 2y=-x-6 y=-1/2x-6 now can you see what the slope is

OpenStudy (anonymous):

is that how u suppsoe to write it out

OpenStudy (anonymous):

you need the slope of it to figure out the slope of the other line. the slope of x+2y=-6 would be -1/2. perpendicular lines have negative reciprocal slopes of each other so the other lines slope would be 2. now use y-y1=m(x-x1) to get the equation of the line

OpenStudy (anonymous):

x-2y=5